# Too Many Frames, Not All Useful: Efficient Strategies for Long-Form Video QA

### v4

@@ 4.1 Experimental Setup
Datasets: Given the training free nature of our framework, we do not utilize any video datasets for training. Datasets are used purely for evaluation. We select three benchmark video visual question answering datasets focused on long-form videos for this purpose: EgoSchema [24], NExT-QA [41], and IntentQA [22]. The first dataset, EgoSchema, is a very long-form video question-answering dataset with 5031 questions. For each question, EgoSchema requires the correct answer to be selected between five given options based on a three-minute-long video clip. The second dataset, NExT-QA, is another rigorously designed video question answering benchmark containing questions that require causal & temporal action reasoning, and common scene comprehension to correctly answer. These questions are further classified as Causal (Cau.), Temporal (Tem.), and Descriptive (Des.) and we evaluate on its validation set containing 4996 questions over 570 videos. The third dataset, IntentQA [22] is based on NExT-QA videos corresponding to temporal and causal reasoning quetions. It consists of 16k multiple-choice questions which are classified as Why?, How? or Before/After (B./A.).
@@
Model Choices: For the HKS module, we use the ResNet-18 [12] for the TSC, CLIP-B/16 [29] for the CKD and GPT-4o for the FKD. We select ResNet-18 and CLIP-B/16 due to their smaller models sizes—0.01B and 0.12B, respectively–which are significantly lighter compared to GPT-4o, whose model size is expected to be on the scale of 100B-1T. This makes them well-suited for filtering dense frames efficiently. In line with previous state-of-the-art work [39, 52, 36], we employ GPT API, especially GPT-4o, for both VLM and LLM. This choice is driven by its cost-effectiveness and lighter computational requirements compared to GPT-4. GPT-4o is used as the VLM for generating captions and as the LLM for answering questions in our framework.
Table 1: Combined Results on EgoSchema [24], NExT-QA [41], and IntentQA [22]. We compare LVNet against prior zero-shot models across three datasets, highlighting different task splits. The models are ordered based on the number of captions processed before answering the question. LVNet achieves state-of-the-art accuracies of 71.7%, 61.1%, and 72.9% on the three datasets, respectively, using just 12 frames compared to the models using the similar number of captions. Models with video-caption pretraining or utilizing significantly more captions than the 12 frames used by LVNet are de-emphasized in grey or downplayed in light green to ensure fairness with image-level pretraining or highlight caption efficiency.
Model EgoSchema NExT-QA IntentQA
Cap. Acc. (%) Cap. Cau. (%) Tem. (%) Des. (%) All (%) Cap. Why? (%) How? (%) B./A. (%) All (%)
Vamos [36] - 48.3 - - - - - - - - - -
IG-VLM [17] - 59.8 - 69.8 63.6 74.7 68.6 - - - - 65.3
VideoLLaMA 2 [5] - 53.3 - - - - - - - - - -
InternVideo2 [38] - 60.2 - - - - - - - - - -
Tarsier [35] - 61.7 - - - - 79.2 - - - - -
VIOLET [10] 5 19.9 - - - - - - - - - -
mPLUG-Owl [46] 5 31.1 - - - - - - - - - -
VideoAgent [37] 8.4 54.1 8.2 72.7 64.5 81.1 71.3 - - - - -
MVU [30] 16 37.6 16 55.4 48.1 64.1 55.2 - - - - -
MoReVQA [25] 30 51.7 16 70.2 64.6 - 69.2 - - - - -
TraveLER [37] - - $\sim$25 70.0 60.5 78.2 68.2 - - - - -
VFC [26] - - 32 45.4 51.6 64.1 51.5 - - - - -
SeViLA [48] 32 22.7 32 61.3 61.5 75.6 63.6 32 - - - 60.9
ProViQ [6] 60 57.1 60 - - - 64.6 - - - - -
VideoTree [40] 63.2 61.1 $\leq$56 75.2 67.0 81.3 73.5 $\leq$56 - - - 66.9
FrozenBiLM [45] 90 26.9 - - - - - - - - - -
LifelongMemory [39] 90 62.1 - - - - - - - - - -
LangRepo [16] 180 41.2 90 64.4 51.4 69.1 60.9 90 62.8 62.4 47.8 59.1
LLoVi [52] 180 50.3 90 69.5 61.0 75.6 67.7 90 68.4 67.4 51.1 64.0
LVNet (ours) 12 61.1 12 75.0 65.5 81.5 72.9 12 75.0 74.4 62.1 71.7
@@ 4.2 Evaluation
We evaluate LVNet on the EgoSchema, NExT-QA, and IntentQA dataset and present our results in Table 1. Models with video-caption pretraining are de-emphasized in grey to ensure fairness with image-level pertaining. Models utilizing significantly more captions than the 12 frames are downplayed in light green to consider caption efficiency. For EgoSchema, we achieve 61.1% on the fullest, the highest among the models utilizing approximately 12 captions. This result outperforms VideoAgent, the next best model using 8.4 captions, by 7%. To ensure a fair comparison, we ablate the number of captions for LVNet in LABEL:tab:ablation:numberofframecaptions, observing that LVNet with 8 captions still outperforms VideoAgent with 8.4 captions by 4.2% in the EgoSchema subset. The superior accuracy of LVNet over other keyframe selectors, such as those used in VideoAgent and TraveLER, is discussed in Section 3.4.
We next evaluate on the NExT-QA[41] dataset. This dataset has a particular focus on both temporal and casual reasoning based question-answer pairs. Our approach achives state-of-the-art performance on this benchmark outperforming prior work among the models utilizing approximately 12 captions. In fact, our LVNet with just 12 frame captions achieves 72.9% overall accuracy, outperforming VideoAgent [37] (71.3% at 8.2 captions) by 1.6%.
We finally evaluate on the IntentQA[22] dataset. Our approach achieves state-of-the-art results on this benchmark, outperforming all prior work, including the de-emphasized models with video-caption pretraining and the downplayed models utilizing significantly more captions than 12 frames. In fact, LVNet shows a substantial improvement (+4.8%) over the next best model, VideoTree [40] (66.9% with 90 captions), achieving 71.7% accuracy with only 12 frame captions.
@@
Qualitative Analysis of the Hierarhical Keyframe Selector:
@@ 4.3 Ablations
Table 2: Ablations on EgoSchema [24]: We evaluate different design decisions of our framework on EgoSchema 500-video subset for zero-shot video VQA.
Visual Templating order Acc. (%)
Temporal 65.2
Hybrid (Confidence + Temporal) 68.2
TSC CKD FKD Acc.
68.2
65.8
64.5
62.6
Frames 6.4 8 8.4 11 12 16
VideoAgent[37] 58.4 - 60.2 57.4 - -
LVNet (ours) - 64.4 - - 68.2 67.8
In this section, we present ablations on key design decisions such as the sorting order in FKD, the number of frames for captions, and the effect of different components in HKS. In all ablations, we use a subset of EgoSchema [24], composed of 500 videos. Additional ablations about Choice of LLM and Effect of Patch Size on Keyword Matching in CKD are in Appendix A.1
@@
In visual templating, prioritizing frames by keyword confidence scores followed by reordering low-confidence frames based on timestamp is more effective than relying solely on temporal order. In this hybrid approach, high-confidence frames are selected from shorter segments by sampling three keyframes per set of eight, grouped by confidence rather than timespan. Conversely, low-confidence keyframes, which are crucial but more visually challenging for keyword matching, are sampled from broader video sections, with three frames per set of eight grouped by timestamp. This hybrid approach outperforms purely temporal ordering by 3%.
@@
We conducted an ablation study on the number of frame captions used in our setup, comparing it to VideoAgent [37], which operates with a similarly small number of captions. Our findings show that LVNet achives its highest accuracy 68.2% with 12 captions, while even its lowest accuracy 64.4% with 8 captions still surpasses VideoAgent’s best accuracy 60.2% with 8.4 captions.
@@
This table demonstrates the impact of incrementally adding the temporal scene clustering (TSC), coarse keyframe detector (CKD), and fine keyframe detector (FKD) modules. Without any of these modules, the model relies on uniform sampling and achieves 62.6%. When TSC is added and 12 frames are selected uniformly, the accuracy increases to 64.5%. Adding both TSC and CKD raises the accuracy to 65.8%. Finally, incorporating all three modules—TSC, CKD, and FKD—into the model, which is LVNet, results in an accuracy of 68.2%. This demonstrates the importance of including all modules in LVNet for optimal performance.
@@ Appendix A.1 Additional Ablations
Table A.3: Ablations on EgoSchema [24]: We evaluate different design decisions of our framework on EgoSchema 500-video subset for zero-shot VQA.
@@
Our results show that GPT-4o outperforms both GPT-4 and GPT-3.5 by 2.8% and 7.2%, respectively. Given that GPT-4o is more cost-effective and lightweight compared to GPT-4, we have selected it as our default LLM.
@@
Since keywords can represent activities spanning the entire image or confined to a small region, we adjust the resolution of the visual feature map output from the spatially aware contrastive image pre-training (CLIP) network [29] to match keywords. Our findings show that higher resolutions lead to better accuracy. In LVNet, we use a 14$\times$14 feature map and determine the confidence level of the keyword by selecting the maximum value between the 14$\times$14 patches and the keyword’s text embedding.

### v5

@@ 4 Experiments
Model EgoSchema NExT-QA IntentQA
Cap. Acc. (%) Cap. Acc. (%) Cap. Acc. (%)
Vamos (Wang et al., 2023) - 48.3 - - - -
IG-VLM (Kim et al., 2024) - 59.8 - 68.6 - 65.3
VideoLLaMA 2 (Cheng et al., 2024) - 53.3 - - - -
InternVideo2 (Wang et al., 2024c) - 60.2 - - - -
Tarsier (Wang et al., 2024a) - 61.7 - 79.2 - -
VIOLET (Fu et al., 2023) 5 19.9 - - - -
mPLUG-Owl (Ye et al., 2023) 5 31.1 - - - -
VideoAgent (Wang et al., 2024b) 8.4 54.1 8.2 71.3 - -
MVU (Ranasinghe et al., 2024) 16 37.6 16 55.2 - -
MoReVQA (Min et al., 2024) 30 51.7 16 69.2 - -
VFC (Momeni et al., 2023) - - 32 51.5 - -
SeViLA (Yu et al., 2024) 32 22.7 32 63.6 32 60.9
ProViQ (Choudhury et al., 2023) 60 57.1 60 64.6 - -
VideoTree (Wang et al., 2024e) 62.4 61.1 (56) 73.5 (56) 66.9
FrozenBiLM (Yang et al., 2022) 90 26.9 - - - -
LifelongMemory (Wang et al., 2024d) 90 62.1 - - - -
TraveLER (Shang et al., 2024) (101) 53.3 (65) 68.2 - -
LangRepo (Kahatapitiya et al., 2024) 180 41.2 90 60.9 90 59.1
LLoVi (Zhang et al., 2023) 180 50.3 90 67.7 90 64.0
LVNet (ours) 12 61.1 12 72.9 12 71.7
Table 2: Long Video Evalation: LVNet achieves state-of-the-art accuracies of 71.7%, 61.1%, and 72.9% on EgoSchema, NExT-QA, and IntentQA datasets respectively using just 12 frames compared to models using a similar number of captions. Models are ordered based on number of captions processed per video. Models with video-caption pretraining or utilizing significantly more captions than 12 frames used by LVNet are de-emphasized in grey or downplayed in light green to ensure fair comparison. Numbers in parentheses () indicate the maximum number of frames used. See Sec. A.2 in appendix for detailed results.
Metric Category VideoAgent VideoTree LVNet
Avg. Frames $\downarrow$ - 24.6 98.0 24.0
Acc.(%) $\uparrow$ Knowledge 52.2 60.7 63.0
Film & TV 42.5 52.5 45.0
Sports Comp. 42.7 48.6 48.0
Artistic Perf. 47.5 51.6 53.0
Life Record 44.7 49.5 45.0
Multilingual 36.6 40.0 53.0
[…]
Average 46.4 53.1 52.4
Table 3: Evaluating on Very Long Videos. Comparison of LVNet (ours) with VideoAgent and VideoTree on the long split of VideoMME. LVNet uses the fewest frames while achieving the highest accuracy in three out of six categories and ranking second in overall performance, slightly below the best score. In the table, bold indicates the best performance, while underlined represents the second-best performance.
@@ 4.1 Experimental Setup
Datasets: Given the training free nature of our framework, we do not utilize any video datasets for training. Datasets are used purely for evaluation. We select three benchmark video visual question answering datasets focused on long-form videos for this purpose: EgoSchema (Mangalam et al., 2023), NExT-QA (Xiao et al., 2021), and IntentQA (Li et al., 2023a). In addition, to further highlight the strength of our approach on longer videos, we include results on VideoMME’s long splitFu et al. (2024). These datasets are public available and can be used freely for academic research. The first dataset, EgoSchema, consists of 5031 questions and each video lasts three-minute and have multiple choice question. The second dataset, NExT-QA, is another rigorously designed video question answering benchmark containing questions that require causal & temporal action reasoning, and common scene comprehension to correctly answer. These questions are further classified as Causal (Cau.), Temporal (Tem.), and Descriptive (Des.) and we evaluate on its validation set containing 4996 questions over 570 videos. The third dataset, IntentQA, is based on NExT-QA videos corresponding to temporal and causal reasoning quetions. It consists of 16k multiple-choice questions which are classified as Why?, How? or Before/After (B./A.). The fourth dataset, VideoMME, consists of very long videos—some up to one hour long, with an average duration of 44 minutes, and provides 900 Q&A.
@@
Model Choices & Hyperparameters: For the HKS module, we use the ResNet-18 (He et al., 2016a) for the TSC, CLIP-B/16 (Ranasinghe et al., 2023) for the CKD and GPT-4o for the FKD. We select ResNet-18 and CLIP-B/16 due to their smaller models sizes—0.01B and 0.12B, respectively–which are significantly lighter compared to GPT-4o, whose model size is expected to be on the scale of 100B-1T. This makes them well-suited for filtering dense frames efficiently. In line with previous state-of-the-art work Wang et al. (2024d); Zhang et al. (2023); Wang et al. (2023), we employ GPT API, especially GPT-4o, for both VLM and LLM. This choice is driven by its cost-effectiveness and lighter computational requirements compared to GPT-4. GPT-4o is used as the VLM for generating captions and as the LLM for answering questions in our framework. We run TSC and CKD on a single NVIDIA RTX A5000, which takes approximately two hours to process 500 questions. We use the default hyperparameters for each vision/language module, as we only perform inference, and set the LLM temperature to 0 to ensure reproducibility. Also, We use single run for our experiments.
@@ 4.2 Evaluation
Model Avg. Frame Captions (Rounded) $\downarrow$
6 7 8 10 11 12 16 32 62 70
VideoAgent 58.4 - 63.2 / 60.2 - 60.8 / 57.4 - - - - -
VideoTree - 61.0 - 61.6 62.2 - - 64.4 66.2 67.0
LVNet (Ours) - - 64.4 - - 68.2 67.8 - - -
Templating Order Acc.
Temporal 65.2
Confidence 67.6
Hybrid (both) 68.2
TSC CKD FKD Acc.
62.6
64.5
65.8
68.2
Table 4: Ablation study on EgoSchema (Mangalam et al., 2023): We evaluate different design decisions of our framework on EgoSchema 500-video subset for zero-shot VQA.
@@
We evaluate LVNet on the EgoSchema, NExT-QA, and IntentQA dataset and present our results in Table 2. Models with video-caption pretraining are de-emphasized in grey to ensure fairness with image-level pertaining. Models utilizing significantly more captions than the 12 frames are downplayed in light green to consider caption efficiency. For EgoSchema, we achieve 61.1% on the fullest, the highest among the models utilizing approximately 12 captions. This result outperforms VideoAgent, the next best model using 8.4 captions, by +7%, is on par with VideoTree while using only 1/5 of the captions, and outperforms TraveLER by +7.8% while utilizing only 12% of the captions.
We next evaluate on the NExT-QA dataset. This dataset has a particular focus on both temporal and casual reasoning based question-answer pairs. Our approach achives state-of-the-art performance on this benchmark outperforming prior work among the models utilizing approximately 12 captions. In fact, our LVNet outperforms VideoAgent by +1.6%.
In the IntentQA dataset. LVNet outperforms all prior work, including the de-emphasized models with video-caption pretraining and the downplayed models utilizing significantly more captions than 12 frames. In fact, LVNet shows a substantial improvement of +4.8% over the next best model, VideoTree, while using only 13% of the captions (12 vs. 90).
Lastly, Table 4 presents the performance of LVNet on VideoMME’s long split, which consists of videos up to one hour long and compare it to other models using keyframe selection methods. Our method (LVNet) demonstrates strong performance while utilizing only 24 frames, significantly fewer than VideoTree’s 98 frames. LVNet outperforms VideoAgent by +6.0% overall and achieves the highest accuracy in three out of six categories: Knowledge, Artistic Performance, and Multilingual. While VideoTree maintains a slight overall lead, LVNet’s ability to achieve comparable accuracy while processing only one-quarter of the frames highlights its efficiency in handling very long videos. To ensure a fair comparison, all models utilize GPT-4o.
@@
Qualitative Analysis of Hierarhical Keyframe Selector:
@@ 4.3 Ablations
In this section, we present ablations on key design decisions such as the sorting order in FKD, the number of frames for captions, and the effect of different components in HKS. In all ablations, we use a subset of EgoSchema (Mangalam et al., 2023), composed of 500 videos. Additional ablations about Choice of LLM and Effect of Patch Size on Keyword Matching in CKD are in Appendix A.1
@@
In visual templating, prioritizing frames by keyword confidence scores followed by reordering low-confidence frames based on timestamp proves more effective than using confidence scores or temporal order alone, as shown in Table LABEL:tab:ablation:visualtemplateorder. In this hybrid approach, high-confidence frames capture short but important segments of the video, while low-confidence keyframes, which are crucial but visually challenging for keyword matching, are temporally ordered to cover broader segments. This hybrid approach outperforms solely temporal ordering and solely confidence-based ordering by +3% and +0.6%, respectively.
@@
We performed an ablation study on the number of frame captions, comparing our approach to VideoAgent (Wang et al., 2024b) and VideoTree (Wang et al., 2024e) under similar low caption settings. As shown in Table LABEL:tab:ablation:numberofframecaptions, LVNet achieves the highest accuracy of 68.2% with 12 captions, outperforming VideoAgent (8.4 frames) and VideoTree (12 frames) by +8% and $\sim$+5.7%, respectively. We compare LVNet with VideoAgent+GPT-4o (8.1 frames) and VideoTree+GPT-4o (69.5 frames, $\times$5.8 more), both using GPT-4o for a fair comparison. We take GPT-4o numbers for VideoAgent and VideoTree from Yang et al. (2024) (more details in appendix A.2) and LVNet outperforms them by +5% and +1.2%, respectively.
@@
Table LABEL:tab:ablation:effectofcomponents demonstrates the impact of incrementally adding the temporal scene clustering (TSC), coarse keyframe detector (CKD), and fine keyframe detector (FKD) modules. Without any of these modules, the model relies on uniform sampling and achieves 62.6%. When TSC is added and 12 frames are selected uniformly, the accuracy increases to 64.5%. Adding both TSC and CKD raises the accuracy to 65.8%. Finally, incorporating all three modules—TSC, CKD, and FKD—into the model, which is LVNet, results in an accuracy of 68.2%. This demonstrates the importance of including all modules in LVNet for optimal performance.
@@ 5 Conclusion
Limitations
Despite the effectiveness of LVNet, as demonstrated by benchmark experiments and comprehensive ablations, our study has certain limitations, which we discuss below.
  • First, we acknowledge that we are unable to evaluate LVNet and other models with all available VLMs or LLMs due to computational constraints and high costs. However, we carefully select GPT-4o, a state-of-the-art LLM, for our main experiments and provide ablation studies comparing various LLMs (e.g. GPT-3.5, GPT-4, and GPT-4o) to other models to ensure a fair performance comparison, as presented in LABEL:tab:ablation:numberofframecaptions and LABEL:tab:ablation:choiceofllm.
  • Our hierarchical keyframe selector consists of three components: TSC, CKD, and FKD. While we demonstrated the effectiveness of each component in LABEL:tab:ablation:effectofcomponents, we did not have the time or resources to develop a unified module that could replace all three. Although this is beyond the scope of this paper, exploring a more efficient implementation that integrates these three modules into a single model would be an interesting direction for future research.
  • Like any LLM-based approach, LVNet is sensitive to prompting. To ensure the transparency, we provide examples of these prompts in Figure 4 and Figure A.7. We also plan to release the code to enable further exploration by other researchers.
  • Finally, we acknowledge that, as our approach is zero-shot, any inherent limitations or biases in the pretrained models may persist in the outputs of LVNet.
@@ Acknowledgements:
Figure A.5: Comparison of Keyframe Selection: Comparison of LVNet and VideoAgent in keyframe selection for video question answering. LVNet refines frames through a multi-stage process (TSC, CKD, FKD) to form a non-uniform keyframe distribution, capturing relevant moments tied to the query. In contrast, VideoAgent relies on uniform sampling and LLM-based frame selection, which fails to focus on crucial keyframes, leading to incorrect predictions. The final keyframe distributions illustrate LVNet’s ability to retrieve meaningful frames directly related to the answer, while VideoAgent selects irrelevant frames.
@@ Appendix A.1 Additional Ablations
Table A.5: Additional ablations experiments on EgoSchema (Mangalam et al., 2023): We evaluate different design decisions of our framework on EgoSchema 500-video subset for zero-shot VQA. Default setting is highlighted.
@@
Table LABEL:tab:ablation:choiceofllm shows that GPT-4o outperforms both GPT-4 and GPT-3.5 by +2.8% and +7.2%, respectively. Given that GPT-4o is more cost-effective and lightweight compared to GPT-4, we have selected it as our default LLM.
@@
Table LABEL:tab:ablation:patchsizeinckd shows the effect of the scales of the patch sizes in the CKD. Since keywords can represent activities spanning the entire image or confined to a small region, we adjust the resolution of the visual feature map output from the spatially aware contrastive image pre-training (CLIP) network (Ranasinghe et al., 2023) to match keywords. Our findings show that higher resolutions lead to better accuracy. In LVNet, we use a 14$\times$14 feature map and determine the confidence level of the keyword by selecting the maximum value between the 14$\times$14 patches and the keyword’s text embedding.
Table A.6: Exact (Unrounded) Frame Caption Counts. These values supplement the rounded numbers in Table LABEL:tab:ablation:numberofframecaptions All models are based on either GPT-4o or GPT-4.
Model Avg. Frame Captions $\downarrow$
6.4 7.1 8 8.1 8.4 9.7 10.7 11 11.3 12 16 32 62.4 69.5
VideoAgent 58.4 - - 63.2 60.2 - 60.8 57.4 - - - - - -
VideoTree - 61.0 - - - 61.6 - - 62.2 62.5 66.2 64.4 66.2 67.0
LVNet (Ours) - - 64.4 - - - - - - 68.2 67.8 - - -
Exact Frame Caption Counts:
For completeness, Table A.6 lists the original (unrounded) frame caption counts and corresponding accuracies for VideoAgentWang et al. (2024b), VideoTreeWang et al. (2024e), and LVNet (ours). These values supplement the rounded numbers presented in Table LABEL:tab:ablation:numberofframecaptions of the main text.
Appendix A.2 Extended results on NExT-QA and IntentQA
We present extended zero-shot evaluation results on NExT-QA in Table A.7, comparing LVNet with prior zero-shot models across different task categories: causal, temporal, and descriptive reasoning. Models are ordered based on the number of captions processed per video, highlighting the trade-offs between caption efficiency and performance.
LVNet achieves state-of-the-art performance with an overall accuracy of 72.9%, outperforming most models while using only 12 captions per video. Notably, it attains 75.0% on causal reasoning, which is the highest among all models evaluated. For temporal reasoning, LVNet achieves 65.5%, remaining competitive despite using significantly fewer captions than models like VideoTree (56 captions) and LangRepo (90 captions). In descriptive reasoning, LVNet reaches 81.5%, matching VideoTree while processing significantly fewer captions.
Compared to VideoAgent, the closest competing model in terms of caption efficiency (8.4 captions), LVNet demonstrates a substantial performance gain across all categories, with a +2.8% improvement in overall accuracy. While models like VideoTree and TraveLER show strong performance, they process significantly more captions (56 and 65, respectively), indicating that LVNet achieves a superior balance between efficiency and accuracy.
We present extended zero-shot evaluation results on IntentQA in Table A.8, comparing LVNet with prior zero-shot models across different reasoning categories: Why?, How?, and B.A. (Before/After). Models are ordered based on the number of captions processed per video, highlighting the balance between caption efficiency and performance.
LVNet achieves an overall accuracy of 71.7%, outperforming all models while using only 12 captions per video. It achieves 75.0% on the Why? category, 74.4% on the How? category, and 62.1% on the B.A. category. Compared to VideoTree, which processes 56 captions and achieves an overall accuracy of 66.9%, LVNet outperforms it by +4.8% while using significantly fewer captions. Similarly, LangRepo and LLoVi, which process 90 captions, achieve overall scores of 59.1% and 64.0%, respectively, further demonstrating LVNet’s caption efficiency.
To ensure fairness, models that utilize video-caption pretraining or process substantially more captions than LVNet are de-emphasized in grey or downplayed in light green in Table A.7 and Table A.8. We adopt the reported GPT-4o results for VideoAgent and VideoTree in Section 4 and LABEL:tab:ablation:numberofframecaptions from the VCA Yang et al. (2024), but do not compare directly against VCA for two reasons: (1) the VCA paper does not provide code or implementation details (e.g., inference speed), making replication infeasible; and (2) its reported results cover only a subset of Egoschema, preventing a fair comparison to our approach on the full-scale EgoSchema, NExT-QA, and IntentQA. Overeall, these clarifications, alongside the results in Table A.7 and Table A.8, underscore LVNet’s effectiveness in achieving high accuracy while maintaining computational efficiency.
Model Cap. Cau. (%) Tem. (%) Des. (%) All (%)
IG-VLM (Kim et al., 2024) - 69.8 63.6 74.7 68.6
Tarsier (Wang et al., 2024a) - - - - 79.2
VideoAgent (Wang et al., 2024b) 8.2 72.7 64.5 81.1 71.3
MVU (Ranasinghe et al., 2024) 16 55.4 48.1 64.1 55.2
MoReVQA (Min et al., 2024) 16 70.2 64.6 - 69.2
VFC (Momeni et al., 2023) 32 45.4 51.6 64.1 51.5
SeViLA (Yu et al., 2024) 32 61.3 61.5 75.6 63.6
VideoTree (Wang et al., 2024e) (56) 75.2 67.0 81.3 73.5
ProViQ (Choudhury et al., 2023) 60 - - - 64.6
TraveLER (Shang et al., 2024) (65) 70.0 60.5 78.2 68.2
LangRepo (Kahatapitiya et al., 2024) 90 64.4 51.4 69.1 60.9
LLoVi (Zhang et al., 2023) 90 69.5 61.0 75.6 67.7
LVNet (ours) 12 75.0 65.5 81.5 72.9
Table A.7: Extended results on NExT-QA (Xiao et al., 2021). We compare LVNet against prior zero-shot models across different reasoning categories: causal, temporal, and descriptive. LVNet achieves an overall accuracy of 72.9% while using only 12 captions per video, demonstrating strong performance across all reasoning types. Notably, it outperforms all models in causal reasoning (75.0%) and matches the best performance in descriptive reasoning (81.5%), despite processing significantly fewer captions than models like VideoTree (56 captions) and TraveLER (65 captions). Models that utilize video-caption pretraining or process substantially more captions than LVNet are de-emphasized in gray or downplayed in light green to ensure fairness in comparison. Numbers in parentheses () indicate the maximum number of frames used.
Model Cap. Why? (%) How? (%) B./A. (%) All (%)
IG-VLM (Kim et al., 2024) - - - - 65.3
SeViLA (Yu et al., 2024) 32 - - - 60.9
VideoTree (Wang et al., 2024e) (56) - - - 66.9
LangRepo (Kahatapitiya et al., 2024) 90 62.8 62.4 47.8 59.1
LLoVi (Zhang et al., 2023) 90 68.4 67.4 51.1 64.0
LVNet (ours) 12 75.0 74.4 62.1 71.7
Table A.8: Extended results on IntentQA (Li et al., 2023a). We compare LVNet against prior zero-shot models across different reasoning categories: Why?, How?, and B.A. (Belief/Action). LVNet achieves an overall accuracy of 71.7%, surpassing all models while using only 12 captions per video. It reaches 75.0% in the Why? category, 74.4% in the How? category, and 62.1% in the B.A. category. Compared to VideoTree, which processes 56 captions and achieves 66.9% accuracy, LVNet outperforms it by +4.8% while using significantly fewer captions. Additionally, LVNet demonstrates superior reasoning-based performance compared to LangRepo (90 captions, 59.1%) and LLoVi (90 captions, 64.0%). Models with video-caption pretraining or utilizing significantly more captions than 12 frames used by LVNet are de-emphasized in grey or downplayed in light green to ensure fairness with image-level pretraining or highlight caption efficiency. Numbers in parentheses () indicate the maximum number of frames used.